OpenStudy (anonymous):

sin theta= 3/5 cos beta = 2sqrt6/5 Find sin(theta+beta). Is the answer 6sqrt6+4/25?

OpenStudy (shamim):

R u able to find out Cos theta=?

OpenStudy (anonymous):

Yes 4/5

OpenStudy (usukidoll):

\[\sin (\theta) = \frac{3}{5}, \cos(\beta) = 2\sqrt{\frac{6}{5}}\] ? I'm confused on what cosine beta should be

OpenStudy (shamim):

Sin beta=?

OpenStudy (anonymous):

1/5

OpenStudy (usukidoll):

oh yeah we need sin beta and cosine beta otherwise the angle sum identity formula isn't going to work

OpenStudy (usukidoll):

\[\sin(\theta + \beta) = \sin(\theta)\cos(\beta)+\cos(\theta)\sin(\beta)\]

OpenStudy (usukidoll):

so far we have sin (theta) = 3/5  so cos (theta) = 4/5

OpenStudy (anonymous):

Yep got that

OpenStudy (usukidoll):

ummm what was the cosine beta?

OpenStudy (usukidoll):

is it \[\cos(\beta) = 2 \frac{\sqrt{6}}{5} ?\] is the square root inside the 6/5? that part is throwing me off

OpenStudy (usukidoll):

\[\sin(\theta + \beta) = \frac{3}{5}\cos(\beta)+\frac{4}{5}\sin(\beta)\] I don't understand what cosine beta is... but I do know that I need another triangle so I can grab sin beta and we add this all together

OpenStudy (usukidoll):

is the 6/5 fraction inside the whole square root or is it just square root 6 / 5 ?

OpenStudy (usukidoll):

\[\cos(\beta) = 2 \frac{\sqrt{6}}{5} ? \] or \[\cos(\beta) = 2 \sqrt{\frac{6}{5}} \]

OpenStudy (usukidoll):

I need clarification on that part... otherwise I might end up with the wrong answer.

OpenStudy (usukidoll):